OpenStudy (anonymous):

Solve for x: e^x^2 - 8 x^2=4

OpenStudy (anonymous):

why? this can't really be done with standard methods

OpenStudy (anonymous):

sure it can..

OpenStudy (anonymous):

differentiating both sides may work

OpenStudy (anonymous):

what is the answer? seems it is in natural log form

OpenStudy (cwrw238):

try this way e^x^2 - 8 x^2=4 e^x^2 = 8 x^2+ 4 taking logs: x^2 = ln (8x^2 + 4) hmm not sure to be honest...

OpenStudy (anonymous):

sqrt(ln 8) ?

OpenStudy (cwrw238):

you might try a graphical method draw graphs of x^2 and the log function on same axes

OpenStudy (anonymous):

been a while since i studied differentiation, but here goes differentiating both sides we get (2x)*e^x^2 -16x =0 if x is not = 0 2*e^x^2 =16 or e^x^2 =8 taking natural log both sides x^2 =ln (8) x=sqrt{ln(8)} whatever this value is...... I may be wrong in my calculations since its been a few years i studied calculus.

OpenStudy (cwrw238):

using graphs i get solution set (-1.86 , 1.86)

OpenStudy (anonymous):

differentiation doesn't tell us anything. finding x for f'(x) = 0 is the location where the slope = 0

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

hmm .....forgot

OpenStudy (cwrw238):

the answers i gave check out but there might be an algebraic method to do this but can't think of one

OpenStudy (anonymous):

there must be an algebraic method,y'all!

OpenStudy (tkhunny):

There is NOT an algebraic method. Even if you simplify it with substitute, \(e^{y} - 8y = 4\), there is NOT algebraic solution. Take a good, hard look at Lambert's W-Function and you will see the similarity and the problem. \(x = \pm\sqrt{ln(8)}\) is the location of the symmetrical absolute minima. The solutions must be outside this range. \(\sqrt{ln(8)}\) might be a good place to start looking, but you should find THEM with slightly greater magnitude. \(\sqrt{ln(8)} = 1.4420268866\) This gives @cwrw238 's graphical solution of +/- 1.86 good credibility. Newton's Method (or just about any other numerical method) MAY find them quickly in this situation. We have to find only one and the other is trivial, by symmetry. Actually, Newton's Method fails miserably when starting at \(x = \sqrt{ln(8)}\). You can get there when starting with x = 1.5, but we already have @cwrw238 's 1.86, so let's start there. Indeed, it converges VERY quickly. \(x_{0} = 1.86\) \(x_{1} = 1.85856070410157\) \(x_{2} = 1.85855496698982\) \(x_{3} = 1.85855496689909\) So, there they are, x = +1.85855496689909 or x = -1.85855496689909 You CANNOT do this with simple algebraic manipulation. There is NOT a nice, closed-form expression, unless you consider the Lambert W-Function amongst the tools that constitute "simple" or "closed-form".

OpenStudy (tkhunny):

Note: When I say there ISN'T a way to do it, this does not mean we're just not clever enough. It does not mean that if we keep trying, we might get it. It can't be done!! That is what it means.

OpenStudy (tkhunny):

The base case is this, \(x = e^{x}\). Can't do it. Try as you might, you will not succeed. Lambert worked with this: \(W(x) = xe^{x}\). Very similar. It can be disappointing to a sharp algebra student to discover that such simple expressions do NOT have lovely algebraic solutions or manipulations that help us solve them. Nevertheless, it is the way it is.

OpenStudy (cwrw238):

yep tkhunny is right

OpenStudy (anonymous):

thanks people!

OpenStudy (cwrw238):

yw